OpenStudy (torii897):

please help Part A: Explain why x = 3 makes 4x − 1 ≤ 11 true but not 4x − 1 < 11. (5 points) Part B: What value from the set {4, 5, 6, 7, 8} makes the equation 8x + 1 = 57 true? Show your work.

OpenStudy (texaschic101):

Part A : 4x - 1 <= 11.....when x = 3 4x - 1 < 11....when x = 3 4(3) - 1 <= 11 4(3) - 1 < 11 12 - 1 <= 11 12 - 1 < 11 11 = 11 (correct) 11 < 11 (false...because 11 is not less then 11) Part B : 8x + 1 = 57 --- subtract 1 from both sides 8x = 57 - 1 8x = 56 -- divide both sides by 8 x = 7.....any other value for x will not work

OpenStudy (texaschic101):

oops..part A....because 11 is not less then 11, it is equal

OpenStudy (torii897):

can I ask u 3 more? @texaschic101